# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

### v1

@@ 1 Introduction
It has become common practice to train application-ready language models in two phases (Radford et al., 2018; Kenton and Toutanova, 2019): first, the model is pre-trained on a very large and general corpus of (unlabeled) text; then further trained (or fine-tuned) on a smaller specific set of examples demonstrating the intended behavior for a particular application, such as instruction following (Ouyang et al., 2022b).
@@
To investigate the performance of SpaRTA, we build adapters with different sparsity levels and test on the GLUE classification tasks (Wang et al., 2019) as well as the IMDB dataset (Maas et al., 2011). SpaRTA is compared to a fully fine-tuned model (Full FT), and PEFT approaches including LoRA. SpaRTA is quite competitive compared to LoRA given that it only modifies a small sparse number of model parameters.
@@ 2 Motivation
Yu et al. (2024) look at the differences between a language model’s parameters before and after fine-tuning, and demonstrate empirically that it is possible to randomly drop (i.e., set to zero) up to 99% of these parameter changes without significantly affecting the model performance. The sparsity in the parameter changes represented by the tensor $\Delta$ is known as $\Delta$-sparsity.
This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters: SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero). This parameter sparsity also helps in reducing overfitting, as pre-trained models typically have enough capacity to learn the often limited amount of (labeled) data used for fine-tuning. There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned, task-specific models into a single one that can perform all tasks simultaneously.
One byproduct of $\Delta$-sparsity is sparse $\Delta$ updates which reduce gradient computations during training, ensure fast merge back of the sparse $\Delta$ into the model for inference, and ultimately same inference cost as the original model (like LoRA). The properties of low memory, low computation needs in training, identical inference time as the original model are all quite desirable for an adaptation technique. SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
2.1 Is low-rank necessary for $\Delta$?
In Appendix A, we show that the weight matrix changes (i.e., $\Delta$) during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b. This indicates that LoRA works, not because of the low-rank constraint particularly, but rather due to the capacity reduction achieved since fine-tuning with LoRA is done with limited training data. Such insight also hints that any constraint which reduces the capacity of the original model could perform competitively, motivating our approach that selects a small number of parameters of the original model to be updated during training.
@@ 3 Related work
The last few years has seen many advances in fine-tuning Large Language Models (LLMs). Perhaps the most well-known method (also most used in practice) is LoRA Hu et al. (2022), which has spurred many variants including DoRA – which adapts only the directions of pre-trained weights Liu et al. (2024), VeRA – which shares low-rank matrices across layers Kopiczko et al. (2024), AdaLoRA – which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al. (2023), and SoRA – which dynamically adjusts the intrinsic rank during the adaptation Ding et al. (2023). This set of methods each have a different structure to the $\Delta$ being optimized with the commonality being the training of some function of low-rank matrices. For example, further expanding on VeRA Kopiczko et al. (2024), the authors propose to learn $\Delta=DBEA$ where $A$ and $B$ are random low-rank matrices and $D$ and $E$ are diagonal matrices to be trained.
Beyond adding structured parameters is the concept of fine-tuning a small subset of the total parameters of the model, i.e., sparse fine-tuning, where one must first decide which subset of parameters to fine-tune (similar to deciding which parameters in each layer to add adapters to in LoRA). Ansell et al. (2024) start with an initial set of parameters and offer procedures to drop (according to parameter magnitude) and grow (according to gradient) the set, i.e., they learn the set of parameters to train. Alternatively, Ma et al. (2024) focus on the sparsity of neuron activations during training by pre-computing neuron importance scores and only including important neurons in computations during training. Ansell et al. (2022) first fine-tune on all parameters, select the parameters that change the most, and then fine-tune again from scratch on the selected parameters. In contrast to these, our proposed SpaRTA method produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters chosen completely at random.
@@ 4 SpaRTA: Adapting a Random Subset of Model Parameters
Suppose the vectorized parameters of the pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) of the model is performed with a labeled dataset characterizing a task. FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm (e.g., Adam from Kingma and Ba (2015)) to maximize the conditional probability of the labels in the training dataset under the model. The FT model is then given by $\theta_{\text{FT}}=\theta_{\text{PT}}+\Delta_{\text{FT}}$ where $\Delta_{\text{FT}}\in\mathbb{R}^{n}$.
This has two drawbacks in terms of memory efficiency. First, the optimizer state can be large, e.g., the state of the Adam optimizer is $4$ times as large as the parameter space as it includes current parameter values, a moving average of gradients (per parameter), etc. Second, storing a new FT model takes as much memory as the PT model which could be an issue when many task-specific models are requested.
SpaRTA proposes to randomly select a small subset of the model parameters to optimize while freezing the rest. The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of parameters selected. This allows our approach to have a drastically lower memory footprint than FT (similar to LoRA) by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
In our SpaRTA’s implementation, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$. The pseudocode for SpaRTA is given in Algorithm 1.
@@
Sample $\phi$: Indices of parameters to be optimized
@@ 5. Update $\Delta_{\phi}$ with the Adam Optimizer
Whereas optimizing LoRA requires computing gradients with respect to the $A$ and $B$’s used to compute $\Delta$ (which constitute additional parameters, independent from $\theta_{\text{PT}}$), SpaRTA requires computing gradients with respect to $\Delta_{\phi}$, the changes over a subset of parameters chosen from $\theta_{\text{PT}}$ and indexed by $\phi$. This nuance is why LoRA can trained with out-of-the-box optimizers, but SpaRTA requires a different implementation. For example, Step 4 in Algorithm 1 must take into account the computational graph so as to not interfere with the forward pass to be computed at the following iteration. Specifically, this unmerge operation must not be recorded in the computational graph.
Generating the index set $\phi$ is done as follows. For each $i\in\{1,\ldots,n\}$, SpaRTA samples $x\sim\mbox{Bernoulli}(m/n)$ and includes $i\in\phi$ if $x=1$. Hence, SpaRTA uses $m$ trainable parameters in expectation. Step 5 also depends on the specific optimizer and settings being used; SpaRTA uses Adam in our implementation. Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes. Thus, SpaRTA does not introduce any additional inference latency when deploying the resulting fine-tuned model.
@@ 5 Memory Usage
In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits for the statistics needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits).
Memory savings appear if and only if
@@
that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required ($s\!>\!0.5$), making less than 50% of the model parameters trainable. For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1.
@@ 6 Experimental Setup
We next detail the framework for conducting the experiments in Section 7. Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) for the desired tasks.
@@ 6.1 Tasks
The main experimental focus is on Natural Language Understanding (NLU) tasks, specifically sequence classification, which involves classifying natural language sequences into a given number of classes. NLU tasks are easier to evaluate than Natural Language Generation (NLG) tasks, as one can use simple metrics (Accuracy, F1, Recall, Precision, etc.) that can avoid the inherent ambiguity of evaluating more general generative tasks. While SpaRTA is also applicable to NLG tasks, they are not used in the following demonstrations due to the challenges associated with their evaluation, typically requiring human judgments as the gold standard measure of performance. Human evaluations can be expensive and time consuming, do not scale easily, and are not as objective as NLU evaluations.
@@ 6.2 Language Models
Starting with available open-weight language models, the goal is to adapt them to perform sequence classification. Two types of trained models are used: base and instruction-tuned models, where the latter have additionally been trained to follow users’ intent when prompted, such as answering questions or responding to instructions. Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma (Team et al., 2024) family, and mistral-7b and mistral-7b-it from the Mistral [^2] We use v0.3 for both models. (Jiang et al., 2023) family, developed by Google and Mistral AI respectively. The "it" suffix refers to instruction-tuned models. They are of particular interest for our experiments as they will show results for models with two different numbers of parameters. All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^3] https://huggingface.co/(HF). Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
@@ 6.3 Task Adaptation
When using a base model for sequence classification, the raw sequences to be classified are input directly into the model. However, when using an instruction model, these sequences are first wrapped into a classification-specific instruction. For example, a possible instruction could be: “Determine if the following sentence has a positive sentiment. Respond Yes or No.”, followed by the sequence itself to be classified.
A (generative) pre-trained Transformer model with a decoder-only architecture has a head that transforms the final hidden state of each token in the input sequence into vocabulary logits. Efficiently adapting this model for sequence classification requires the swap of this head for a sequence classification head, which uses only the final hidden state of the last token in the input sequence $h\in\mathbb{R}^{d}$ to do the classification. This reduces the parameter size of the head, which is just a linear layer with no bias applied to a token’s final hidden state, from a weight matrix $W\in\mathbb{R}^{v\times d}$ to $W\in\mathbb{R}^{c\times d}$, where $d$ is the dimension of the model’s hidden states (e.g., 2,048 and 4,096 for the gemma-2b and mistral-7b models respectively), $v$ is the number of tokens in the vocabulary (e.g., 256,000 for gemma-2b or 32,768 for mistral-7b), and $c$ is the number of classes in the individual task (e.g. $2$ to $3$ in the experiments that follow). With this small change, the model outputs classification probabilities for each whole input sequence through
@@
The classification head of our base pre-trained models (i.e., gemma-2b and mistral-7b) is initialized with random weights, sampled from a normal distribution $\mathcal{N}(0,\sigma^{2})$, with $\sigma$ chosen such that the model at initialization predicts either class with equal probability. This is achieved with $\sigma=0.06/\sqrt{d}$. With this, the loss starts at approximately $-\log(1/2)\approx 0.693$ and the accuracy at $\approx 50\%$ for a balanced evaluation dataset.
Regarding instruction-tuned models (i.e., gemma-2b-it and mistral-7b-it), the original vocabulary weights are rather reused when initializing their classification heads. To do so requires to first identify the tokens in the vocabulary that the model is expected to use for classification following the instruction. For example, these could be the tokens associated with a “Yes” or “No” response by the model. The embeddings in the original model (head) associated with those classification tokens are extracted and used to initialize the classification head. While many models tie their vocabulary heads to their tokens embedding matrices, these new classification heads are never tied to the model’s input embedding matrix.
@@ 7 Experimental Results
Efficacy of SpaRTA is demonstrated empirically by comparing it to the following baselines:
  • Full parameter Fine-Tuning (Full FT): all model parameters are optimized.
  • LoRA: Only the additional matrices representing low-rank adapters are optimized.
  • Head: Only the classification head is optimized; all other parameters remain frozen.
This enables the exploration of the performance of SpaRTA on a range of sparsity levels, varying the number of trainable parameters. A complete set of results can be found in Appendix D accompanied by a detailed description of the training setup in Appendices C and E.
7.1 Datasets
Each sequence classification task considered in our experiments is given by a dataset of labeled examples. Table 2 summarizes these datasets and their splits for training, development, and testing for both the GLUE classification tasks (Wang et al., 2019) and IMDB (Maas et al., 2011). See Appendix B for more detailed descriptions.

### v2

@@ 1 Introduction
It has become common practice to train application-ready language models in two phases (Radford et al., 2018; Kenton and Toutanova, 2019): first, the model is pre-trained on a very large and general corpus of (unlabeled) text; then further trained (or fine-tuned) on a smaller specific set of examples demonstrating the intended behavior for a particular application, such as instruction following (Ouyang et al., 2022).
@@
To investigate the performance of SpaRTA, we build adapters with different sparsity levels and evaluate them on a wide range of natural language understanding benchmarks. SpaRTA is compared to other PEFT approaches including LoRA and DoRA Liu et al. (2024), and found to be quite competitive compared to these methods given that it only modifies a small sparse number of model parameters.
@@ 2 Motivation
Yu et al. (2024) look at the differences between a language model’s parameters before and after fine-tuning, and demonstrate empirically that it is possible to randomly drop (i.e., set to zero) up to 99% of these parameter changes, represented by $\Delta$, without significantly affecting model performance.
This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters: SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero). This $\Delta$-sparsity also helps in reducing overfitting, as pre-trained models typically have more than enough capacity to learn the often limited amount of (labeled) data used for fine-tuning. There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned models.
SpaRTA, like LoRA, reduces the number of gradient computations during training (when compared with full Fine-Tuning), and ultimately has the same inference cost as the original model (after merging back the sparse $\Delta$ into the model). The properties of low memory, low computation needs in training, identical inference time as the original model are all quite desirable for an adaptation technique. SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
2.1 Is low-rank adaption necessary?
In Appendix A, we show that the changes (i.e., $\Delta$) in weight matrix during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b. This indicates that LoRA works, not particularly because of its low-rank constraint, but rather due to the capacity reduction achieved since LoRA fine-tuning with is typically done on limited training data. Such insight also hints that any constraint which reduces the capacity of the original model could perform competitively, motivating our approach that selects a small number of parameters of the original model to be updated during training.
@@ 3 Related work
The last few years has seen many advances in fine-tuning Large Language Models (LLMs). Perhaps the most well-known and used method in practice is LoRA Hu et al. (2022), which has spurred many variants including: DoRA, which adapts only the directions of pre-trained weights Liu et al. (2024); VeRA, which shares low-rank matrices across layers Kopiczko et al. (2024); AdaLoRA, which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al. (2023); and SoRA, which dynamically adjusts the intrinsic rank during adaptation Ding et al. (2023). Each of these methods has a different structure to the $\Delta$ being optimized, with the commonality being the training of some function of low-rank matrices. AdaLoRa and SoRA are also dynamic methods, which adjust the function during adaptation.
Beyond adding structured parameters is the concept of fine-tuning a small subset of the total parameters of the model, i.e., sparse fine-tuning, where one must first decide which subset of parameters to fine-tune (similar to deciding which parameters in each layer to add adapters to in LoRA). Ansell et al. (2024) start with an initial set of parameters and offer procedures to drop (according to parameter magnitude) and grow (according to gradient) the set, i.e., they learn the set of parameters to train. Alternatively, Ma et al. (2024) focus on the sparsity of neuron activations during training by pre-computing neuron importance scores and only including important neurons in computations during training. Ansell et al. (2022) first fine-tune on all parameters, select the parameters that change the most, and then fine-tune again from scratch on the selected parameters. Deng et al. (2025) proposes improvements via a modified pruning strategy to the random dropping and rescaling method of Yu et al. (2024) which motivates this work. In contrast to these more complex, sometimes dynamic approaches, our proposed SpaRTA method produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters chosen completely at random.
@@ 4 SpaRTA: Adapting a Random Subset of Model Parameters
Suppose the vectorized parameters of the pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) of the model is performed with a labeled dataset characterizing a task. FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm, e.g., Adam from Kingma and Ba (2015), to maximize the conditional probability of the labels in the training dataset under the model. The FT model is then given by $\theta_{\text{FT}}=\theta_{\text{PT}}+\Delta_{\text{FT}}$ where $\Delta_{\text{FT}}\in\mathbb{R}^{n}$.
This has two drawbacks in terms of memory efficiency. First, the optimizer state can be large, e.g., the state of the Adam optimizer is $4$ times as large as the parameter space as it includes current parameter values, their gradients as well as per parameter statistics of those gradients. Second, storing a new FT model takes as much memory as the PT model which could be an issue when many task-specific models are requested.
SpaRTA proposes to randomly select a small subset of the model parameters to optimize while freezing the rest. The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of parameters selected. This allows our approach to have a drastically lower memory footprint than FT by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
In our SpaRTA’s implementation [^1] https://github.com/IBM/SpaRTA, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$. The pseudocode for SpaRTA is given in Algorithm 1.
@@
Sample $\phi$: Indices of (scalar) parameters to be optimized
@@ 5. Update $\Delta_{\phi}$ with the Adam Optimizer
Whereas optimizing LoRA requires computing gradients with respect to the $A$ and $B$’s used to compute $\Delta$ (which constitute additional parameters, independent from $\theta_{\text{PT}}$), SpaRTA requires computing gradients with respect to $\Delta_{\phi}$, the changes over a subset of parameters chosen from $\theta_{\text{PT}}$ and indexed by $\phi$.
Generating the index set $\phi$ in SpaRTA is done by sampling from a Bernoulli independently and including each scalar parameter in $\phi$ with probability $m/n$. Hence, SpaRTA uses $m$ trainable parameters out of $n$ in expectation. Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes. Thus, SpaRTA does not introduce any additional inference latency when deploying the resulting fine-tuned model.
@@ 5 Memory Usage
In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits for the statistics needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits). Memory savings appear then if and only if
@@
that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required ($s\!>\!0.5$), making less than 50% of the model parameters trainable. For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1. Additional savings are possible regarding storing indices given a fixed random number generator; in this case, the random path of selected parameters to train can be derived given a fixed seed.
@@ 6 Experimental Setup
We next detail our experimental framework. Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) to the tasks.
@@ 6.1 Tasks
The main experimental focus is on Natural Language Understanding (NLU) tasks, specifically sequence classification, which involves classifying natural language sequences into a given number of classes. NLU tasks are easier to evaluate than Natural Language Generation (NLG) tasks, as one can use simple metrics (Accuracy, F1-scores, etc.) that can avoid the inherent ambiguity of evaluating more general generative tasks. While SpaRTA is also applicable to NLG tasks, they are not used in the following demonstrations due to the challenges associated with their evaluation, typically requiring human judgments as the gold standard measure of performance. Human evaluations can be expensive and time consuming, do not scale easily, and are not as objective as NLU evaluations.
@@ 6.2 Language Models
Starting with available open-weight language models, the goal is to adapt them to perform a selected NLU task. Two types of trained models are used: base and instruction-tuned models, where the latter have additionally been trained to follow users’ intent when prompted, such as answering questions or responding to instructions. Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma family (Team et al., 2024), and mistral-7b and mistral-7b-it from the Mistral [^3] We use v0.3 for both models. family (Jiang et al., 2023). The "it" suffix refers to instruction-tuned models. They are of particular interest for our experiments as they will show results for models with two different numbers of parameters. All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^4] https://huggingface.co/. Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
@@ 6.3 Task Adaptation
When using a base model for a sequence classification task, the raw text to be classified is input directly into the model as a sequence of tokens (with some special token if necessary to deal with structured inputs). However, when using an instruction model, these sequences are first wrapped into a classification-specific instruction. For example, a possible instruction could be: “Determine if the following sentence has a positive sentiment. Respond Yes or No.”, followed by the sequence itself to be classified.
A (generative) pre-trained Transformer model with a decoder-only architecture has a head that transforms the final hidden state of each token in the input sequence into vocabulary logits. Efficiently adapting this model for sequence classification requires the swap of this head for a sequence classification head, which uses only the final hidden state of the last token in the input sequence $h\in\mathbb{R}^{d}$ to do the classification. This reduces the parameter size of the head, which is just a linear layer applied to $h$, from a weight matrix $W\in\mathbb{R}^{v\times d}$ to $W\in\mathbb{R}^{c\times d}$, where $d$ is the dimension of the model’s hidden states (e.g., 2,048 and 4,096 for the gemma-2b and mistral-7b models respectively), $v$ is the number of tokens in the vocabulary (e.g., 256,000 for gemma-2b or 32,768 for mistral-7b), and $c$ is the number of classes (e.g. $2$ to $4$ in the experiments that follow). With this small change, the model outputs classification probabilities for each input sequence through
@@
The classification heads of our base pre-trained models (i.e., gemma-2b and mistral-7b) are initialized with random weights. The weights of the the original vocabulary heads of instruction-tuned models (i.e., gemma-2b-it and mistral-7b-it) are rather reused when initializing their classification heads. To do so requires to first identify the tokens in the vocabulary that the model is expected to use for classification following the instruction. For example, these could be the tokens associated with a “Yes” or “No” response. The embeddings in the original model (head) associated with those classification tokens are extracted and used to initialize the classification head. While many models tie their vocabulary heads to their tokens embedding matrices, these new classification heads are never tied to the model’s input embedding matrix.
@@ 7 Experimental Results
Efficacy of SpaRTA is demonstrated empirically by comparing it to two baselines, LoRA (only the additional matrices representing low-rank adapters are optimized) and DoRA (magnitude is optimized in addition to low-rank adapters). We also explore the performance of SpaRTA on a range of sparsity levels, varying the number of trainable parameters.
@@
We consider several NLU benchmarks, including IMDB (Maas et al., 2011), GLUE (Wang et al., 2019), BoolQ (Clark et al., 2019) and MMLU (Hendrycks et al., 2020). See Appendix B for detailed descriptions. Table 2 summarizes these datasets and our splits for training, development, and testing. A detailed description of the training setup can be found in Appendices C and E. All results are averaged over 3 random seeds.